In a parallelogram, angle A = 3x + 5 & angle B = 10(x-2). Find all the angles of the parallelogram. 

Grade:8

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The sum of consecutive angles of a parallelogram is equal to 180 degrees.
Hence, we get:
3x+5+10(x-2)=180^{\circ} \Rightarrow 13x+5-20=180^{\circ}\rightarrow \Rightarrow 13x=180^{\circ}+15^{\circ}=195^{\circ} \Rightarrow x=\frac{195}{13}=15^{\circ}Hence, the required angles are:
A=3x+5=45^{\circ}+5^{\circ}=50^{\circ}; B = 10(x-2)=10(15-2)10\times 13=130^{\circ}Regards